# Runtime Safety through Adaptive Shielding: From Hidden Parameter Inference to Provable Guarantees

### v1

@@ Appendix G Why Not $Q_{C}$, But $Q_{\text{safe}}$?
To effectively guide the policy towards safe behavior, we propose a safety-regularized objective $Q_{\text{safe}}$. A natural alternative is to augment the reward value function $Q_{R}$ with the cost value function $Q_{C}$, which estimates the expected cost of violating safety constraints, forming $Q_{\text{aug}}=Q_{R}-\alpha Q_{C}$. However, this formulation can be transformed into Lagrangian-based safe RL methods, optimizing policies with $Q_{R}-\lambda Q_{C}$, where $\lambda$ is a Lagrangian multiplier dynamically adjusted during training. In particular, the Lagrangian multiplier $\lambda$ is updated using a learning rate $lr$. A higher learning rate accelerates the increase of $\lambda$, assigning stronger penalties on the policy for cost violations. $\lambda$ is adjusted by $Q_{C}\times lr$; larger $Q_{C}$ or learning rate values lead to faster $\lambda$ growth, which increases the penalty term in the optimization objective ($Q_{R}-\lambda Q_{C}$).
@@
However, Lagrangian-based methods are sensitive to the choice of $lr$, often leading to unstable optimization or suboptimal safety-performance trade-offs as shown in Figures 14 and 15. This is because the value of $Q_{C}$ is highly environment-dependent, varying with the magnitude of costs and the dynamics induced by hidden parameters. In contrast, our safety-regularized objective $Q_{\text{safe}}$ incorporates a normalized term, constrained to $(0,1]$. This normalization simplifies controlling the safety bonus by ensuring it remains bounded. Figures 6 and 7 demonstrate that $Q_{\text{safe}}$ maintains consistent performance regardless of $Q_{C}$’s scale, achieving reliable policy optimization in hidden-parameter settings.
Figure 14: Training Curves for Safe Gymnasium with Car Robot. (Top) Average episodic return versus training steps. (Bottom) Average episodic cost rate versus training steps.
Figure 15: Training Curves for Safe Gymnasium with Point Robot. (Top) Average episodic return versus training steps. (Bottom) Average episodic cost rate versus training steps.
Appendix H Cost Functions
@@
$C(s_{t},a_{t},s_{t+1})=\mathbb{I}\left\{\nu(h(s_{t+1}),E_{t+1})\leq 0\right\},$
@@
  • $h:S\to\mathbb{R}^{n_{1}}$ extracts agent-centered safety features from the next state $s_{t+1}$,
@@
Task $h(s)$ $E_{t}$ $\nu\!\bigl{(}h,E\bigr{)}$
Collision avoidance $\operatorname{pos}(s)\!\in\!\mathbb{R}^{3}$$\{X_{i}\}_{i=1}^{M}\!\subset\!\mathbb{R}^{3}$ $\min_{i}\|h-X_{i}\|_{2}-d_{\text{safe}}$
Safety-region compliance $\operatorname{pos}(s)\!\in\!\mathbb{R}^{2}$ $\mathcal{S}_{\mathrm{safe}}\!\subset\!\mathbb{R}^{2}$ $\operatorname{dist}\!\bigl{(}h,\mathbb{R}^{2}\!\setminus\!\mathcal{S}_{\mathrm{safe}}\bigr{)}-\varepsilon$
@@
Thus $C_{d}=1$ whenever the robot violates the distance constraint, encouraging policies that keep a safe distance to obstacles. I
@@
Appendix I Average Cost Minimization and Cost Value Function
Our problem formulation targets minimizing the average cost per time step, distinct from the cumulative discounted cost over an infinite horizon typically addressed by Lagrangian-based methods like TRPO-Lag and PPO-Lag. The connection between cumulative discounted cost and average cost is well-established [Puterman, 2014]:
$\lim_{\gamma\rightarrow 1^{-}}(1-\gamma)V_{C}^{\pi}(s_{0},\theta_{0})=\phi^{\pi}(s_{0},\theta_{0}),$
where $V_{C}^{\pi}(s_{0},\theta_{0})$ denotes the value function for the cost under policy $\pi$ starting from state $s$, parameter $\theta$, and $\phi^{\pi}(s,\theta)=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\theta}}\left[\sum_{t=0}^{H-1}C_{d}(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\theta_{0}\right]$ represents the expected average cost for parameter $\theta$. Thus, by setting the constraint limit to zero in Lagrangian optimization, we ensure the average cost approaches zero, aligning with our safety objectives.
Appendix J Experimental Details
@@ Appendix I Average Cost Minimization and Cost Value Function
For out-of-distribution (OOD) evaluation, we modify Safety Gymnasium task environments—Goal, Button, and Push—by adding two additional hazard locations to increase complexity. For Circle task, we reduce the safety region’s x-limit from 1.125 to 1.0 to further challenge the agent’s safety mechanisms.
To introduce varying hidden parameters, each episode independently samples density, damping, and gravity multipliers by randomly selecting one of two intervals, $[0.1,0.25]$ or $[1.75,2.5]$, with equal probability and uniformly sampling a value from the chosen interval, ensuring diverse environmental conditions.
@@
For runtime execution comparisons, we standardize the hardware to ensure fairness. Experiments are run on a server with 256 GB of memory, dual AMD CPUs (56 cores, 224 threads), and Nvidia RTX A4500 GPUs (20 GB RAM each).

### v2

@@ Appendix I Additional Experiments: Safe Velocity Control in HalfCheetah
To effectively guide the policy toward safe behavior, we propose a safety-regularized optimization enhanced with $Q_{\text{safe}}$. A natural alternative is to augment the reward value function $Q_{R}$ with the cost value function $Q_{C}$, which estimates the expected cost of violating safety constraints, forming $Q_{\text{aug}}=Q_{R}-\alpha Q_{C}$. However, this formulation can be transformed into Lagrangian-based safe RL methods, optimizing policies with $Q_{R}-\lambda Q_{C}$, where $\lambda$ is a Lagrangian multiplier dynamically adjusted during training. In particular, the Lagrangian multiplier $\lambda$ is updated using a learning rate $lr$. A higher learning rate accelerates the increase of $\lambda$, assigning stronger penalties on the policy for cost violations. $\lambda$ is adjusted by $Q_{C}\times lr$; larger $Q_{C}$ or learning rate values lead to faster $\lambda$ growth, which increases the penalty term in the optimization objective ($Q_{R}-\lambda Q_{C}$).
@@
However, Primal-Dual methods such as PPOLag or RCPO are sensitive to the choice of $lr$, often leading to unstable optimization or suboptimal safety-performance trade-offs as shown in Figures 10 and 11. This is because the value of $Q_{C}$ is highly environment-dependent, varying with the magnitude of costs and the dynamics induced by hidden parameters. In contrast, our safety-regularized objective $Q_{\text{safe}}$ incorporates a normalized term, constrained to (1,0](-1,0]. This normalization simplifies controlling the safety bonus by ensuring it remains bounded.
Figure 10: Training curves for RCPO algorithm under varying Lagrangian learning rates. The plots illustrate significant performance variations depending on the learning rate. For our main comparisons, we selected a learning rate of 0.035, which achieves the best trade-off between reward maximization and constraint satisfaction.
Figure 11: Training curves for PPOLag algorithm under varying Lagrangian learning rates. The plots illustrate significant performance variations depending on the learning rate. For our main comparisons, we selected a learning rate of 0.035, which achieves the best trade-off between reward maximization and constraint satisfaction.
Appendix K Fix Parameters vs. Varying Hidden Parameters
Figure 12: Varying hidden parameters pose significant challenges for safe RL algorithms, even when hidden parameter values are provided as input. The first column plot aggregates results across six algorithms: vanilla Lagrangian methods (RCPO and PPO-Lag), trust region/projection methods (CPO and CUP), and safety augmentation techniques (PPO-Saute and TRPO-Saute). Results with varying hidden parameters are shown as solid lines, while those with fixed parameters are depicted as dotted lines. Columns 1, 2, and 3 present comparative results across these algorithm groups.
We evaluate the algorithms under two distinct experimental settings to test performance difference when environment dynamics shifts:
  • Fixed Parameters: In this setting, each algorithm is trained and evaluated in an environment with a single, constant set of parameters $\phi$ (gravity, damping, density, mass, friction) for the entire duration of training.
  • Varying Hidden Parameters: In contrast, for this setting, the underlying physical parameters $\phi$ of the environment such as gravity, damping, mass, inertia, and friction are randomized at the start of each new episode. To demonstrate the challenge of adapting to varying hidden parameters, we explicitly inform the algorithms of these changes via their input. For example, if gravity is halved from 9.8 to 4.9, a factor of 0.5 is provided as input to the policy.
As demonstrated in Figure 12, we observe a noticeable degradation in the performance of all algorithms under the varying parameter setting, despite being explicitly informed of the magnitude of the changes. More precisely, for Point environment, the total aggregated return across training differs significantly: 25.87 for fixed parameters versus 16.49 for varying parameters. Similarly, for the cost, the values are 18.0 for fixed parameters and 38.41 for varying parameters, reflecting more than double the total cost violations during training.
Appendix L Cost Functions
@@
$C(s_{t},a_{t},s_{t+1})=\mathbb{I}\left\{\nu(e(s_{t+1}),E_{t+1})\leq 0\right\},$
@@
  • $e:S\to\mathbb{R}^{n_{1}}$ extracts agent-centered safety features from the next state $s_{t+1}$,
@@
Task $e(s)$ $E_{t}$ $\nu\!\bigl(e(s),E\bigr)$
Collision avoidance $\operatorname{pos}(s)\!\in\!\mathbb{R}^{3}$ $\{X_{i}\}_{i=1}^{M}\!\subset\!\mathbb{R}^{3}$ $\min_{i}\|e(s)-X_{i}\|_{2}-d_{\text{safe}}$
Safety-region compliance $\operatorname{pos}(s)\!\in\!\mathbb{R}^{2}$ $\mathcal{S}_{\mathrm{safe}}\!\subset\!\mathbb{R}^{2}$ $\operatorname{dist}\!\bigl(e(s),\mathbb{R}^{2}\!\setminus\!\mathcal{S}_{\mathrm{safe}}\bigr)-\varepsilon$
@@
Thus $C_{d}=1$ whenever the robot violates the distance constraint, encouraging policies that keep a safe distance to obstacles.
@@
Appendix M Average Cost Minimization and Cost Value Function
Our problem formulation targets minimizing the average cost per time step, distinct from the cumulative discounted cost over an infinite horizon typically addressed by Lagrangian-based methods like TRPO-Lag and PPO-Lag. The connection between cumulative discounted cost and average cost is well-established (puterman2014mdp):
$\lim_{\gamma\rightarrow 1^{-}}(1-\gamma)V_{C}^{\pi}(s_{0},\phi_{0})=\xi^{\pi}(s_{0},\phi_{0}),$
where $V_{C}^{\pi}(s_{0},\phi_{0})$ denotes the value function for the cost under policy $\pi$ starting from state $s$, parameter $\phi$, and $\xi^{\pi}(s,\phi)=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\phi}}\left[\sum_{t=0}^{H-1}C_{d}(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\phi\right]$ represents the expected average cost for parameter $\phi$.
Appendix N Experimental Details
@@
For out-of-distribution (OOD) evaluation, we modify Safety Gymnasium task environments: Goal, Button, and Push, by adding two additional hazard locations to increase complexity. For Circle, we keep the same layout since the wall already blokcs four sides.
To introduce varying hidden parameters, each episode independently samples gravity, damping, mass, inertia, friction multipliers by randomly selecting one of two intervals, $[0.15,0.3]$ or $[1.7,2.5]$, with equal probability and uniformly sampling a value from the chosen interval, ensuring diverse environmental conditions. For Circle task, all settings remain the same except for damping, which is sampled from $[1.7,2.5]$. This adjustment addresses instability in MuJoCo simulator when combined with Circle task, where agents are expected to learn circling behavior. Lower damping factors render the simulator unstable, necessitating this range.